OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 2x + 6y = 18 3x − 2y = 6

OpenStudy (anonymous):

a. −3y + 9 b. −3y − 9 c. 3y − 9 d. 3y + 9

OpenStudy (anonymous):

a) -3y+9 Simplify your equation by isolating the x \[2x+6y=18 \rightarrow 2x=-6y+18 \rightarrow x=-\frac{ 6 }{ 2 } y+\frac{ 18 }{ 2 }\rightarrow x=-3y+9\]

OpenStudy (anonymous):

thanks :D